# Learning in Structured Stackelberg Games

### v2

@@ 3.3 Stackelberg-Littlestone Dimension
Definition 3.6 (Shattered Stackelberg-Littlestone Tree).
[…]
  • (Consistency) $\pazocal{H}$ is consistent with every path in $\pazocal{T}_{d}$ i.e., for any $s\in S_{d}$, there exists $h\in\pazocal{H}$ such that for all $i\in\{\varnothing,1,\dots,d-1\}$ the label of edge $(\mathbf{z}_{s[:i]},\mathbf{z}_{s[:i+1]})$ is $h(\mathbf{z}_{s[:i]})$.
  • (Max-branch property) $\pazocal{T}_{d}$ satisfies the max-branch property, i.e., every node, $\mathbf{z}_{s}$, has children $\mathbf{z}_{si}$ for all $i\in[K]$ where $\pazocal{H}$ is consistent with path $si$.
  • (Weights) each node, $\mathbf{z}_{s}\in\mathscr{T}_{d}$ has a weight, $\rho_{s}$ defined as follows: if $\text{len}(s)=d+1$ then $\rho_{s}=0$, else
    $\rho_{s}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{sj\in S_{d}}\biggl{[}r(\mathbf{z},\mathbf{x},f^{(j)})+\rho_{sj}\biggr{]},$
    where $r(\mathbf{z},\mathbf{x},j):=u(\mathbf{z},\pi_{h^{*}}(\mathbf{z}),b_{f^{(j)}}(\mathbf{z},\pi_{h^{*}}(\mathbf{z})))-u(\mathbf{z},\mathbf{x},b_{j}(\mathbf{z},\mathbf{x}))$ is the leader’s instantaneous regret, given context $\mathbf{z}$, mixed strategy $\mathbf{x}$, and follower type $j$ [^4] The infimum might not be achieved due to the followers’ tie breaking rule. We refer the reader to Section A.1 for further details.
Definition 3.7 (Stackelberg-Littlestone dimension).
The Stackelberg-Littlestone dimension of a hypothesis class $\mathscr{H}$ with respect to a Stackelberg game, $\pazocal{G}$, denoted as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$, is the supremum root node weight value among all finite node-weighted trees that are shattered by hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$.
Note that taking the supremum is necessary since it is possible that no finite-depth tree achieves the optimal value. We begin by showing that no deterministic algorithm can achieve regret lower than the Stackelberg-Littlestone dimension.
Theorem 3.8.
For any hypothesis class, $\pazocal{H}$, any Stackelberg game $\pazocal{G}$, and any deterministic algorithm, there exists an adversarial sequence of contexts and follower types such that the learner suffers at least $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})-\epsilon$ regret, for any $\epsilon>0$.
@@
We prove by induction on the depth $d$ of shattered trees that across $d$ timesteps the learner can be forced to incur regret at least as much as the weight of the tree’s root node. Since the Stackelberg-Littlestone dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim. In case the supremum in the definition of $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ is not achieved by any finite tree, there will always be an additive error of $\epsilon$ that will get arbitrarily small as $d\to\infty$.
(Base Case) We consider a shattered tree of depth 111. Then
@@
(Inductive step) We assume that for any tree of depth $d$ that is shattered by $\pazocal{H}$ under $\pazocal{G}$, the weight of the root node lower bounds the learner’s cumulative regret across $d$ timesteps. We now consider a tree $\mathscr{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$. We will prove that for such a tree, the adversary can guarantee regret at least $\rho_{\varnothing}$, where $\mathbf{z}_{\varnothing}$ is the root node of this tree. By Definition˜3.6, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl{[}r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{i}\biggr{]}$. Let $\mathbf{x}_{*}\in\Delta(\pazocal{A})$ and $i_{*}\in[K]$ be the respective arginf and argmax values in the above expression.
@@
$\implies r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ $\leq\rho_{\tilde{i}}-\rho_{i_{*}}$
We denote $\ell:=r(\mathbf{z}_{\varnothing},\mathbf{x}_{*},f^{(i_{*})})-r(\mathbf{z}_{\varnothing},\tilde{\mathbf{x}},f^{(\tilde{i})})$ and by definition of $\tilde{\mathbf{x}}$ we know that $\ell>0$. We get that $\rho_{\tilde{i}}\geq\rho_{i_{*}}+\ell$. Now by the boundedness of the utility function, we know that if $\rho_{i_{*}}+\ell\geq d$, then the above inequality leads to a contradiction (the learner cannot incur strictly more than $d$ regret over $d$ timesteps).
@@
3.4 Stackelberg Standard Optimal Algorithm
In this section, we present an algorithm that achieves instance-optimal regret that is equal to the Stackelberg-Littlestone dimension. We call Algorithm˜1 the Stackelberg Standard Optimal Algorithm (SSOA), which maintains a similar approach to the Standard Optimal Algorithm (SOA) from online multiclass prediction. Specifically, it keeps track of the subset of hypotheses that are consistent with the history so far (˜8) and plays the mixed strategy that minimizes the worst-case possible instantaneous regret “regularized” by the Stackelberg-Littlestone dimension of the induced hypothesis class (˜6). The latter works as a proxy to the difficulty of the learning task of the leader that would remain after the current round is over.
Algorithm 1 Stackelberg Standard Optimal Algorithm (SSOA)
1:input: time horizon, $T$, Stackelberg game $\pazocal{G}$, and hypothesis class $\pazocal{H}\subseteq[K]^{\pazocal{Z}}$
2:Set $V_{1}=\pazocal{H}$
3:for $t\in[T]$ do
4:  Observe $\mathbf{z}_{t}$
5:  Let $u_{*}^{(i,\mathbf{z}_{t})}:=\sup_{\mathbf{x}\in\Delta(\pazocal{A})}u(\mathbf{z}_{t},\mathbf{x},b_{f^{(i)}}(\mathbf{z}_{t},\mathbf{x}))$ for all $i\in V_{t}(\mathbf{z}_{t})$
6:  Play $\mathbf{x}_{t}=\arg\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{j\in V_{t}(\mathbf{z}_{t})}\left[u_{*}^{(j,\mathbf{z}_{t})}-u(\mathbf{z}_{t},\mathbf{x},b_{f^{(j)}}(\mathbf{z}_{t},\mathbf{x}))+\textnormal{SLdim}_{\pazocal{G}}\left(V_{t}^{(\mathbf{z}_{t}\to j)}\right)\right]$
7:  Observe $f_{t}$
8:  Set $V_{t+1}=V_{t}^{(\mathbf{z}_{t}\to f_{t})}$
9:end
In Theorem˜3.9, we prove the optimality of our algorithm. Our proof follows a high-level idea similar to that of the Standard Optimal Algorithm (SOA) from online multiclass learning: with each mistake, the learner’s task becomes easier, as the Littlestone dimension decreases. Similarly here we show that when the learner incurs positive instantaneous regret, the Stackelberg Littlestone dimension decreases by at least this amount.
Theorem 3.9.
For any Stackelberg game, $\pazocal{G}$, and every hypothesis class, $\pazocal{H}$, the regret of Algorithm˜1 is at most $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$.
@@ 3.4 Stackelberg Standard Optimal Algorithm
We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the Stackelberg-Littlestone dimension must have decreased by at least $r_{t}$. In other words, $\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t})-r_{t}$, where $V_{t}$ is the set of hypotheses consistent with the history up to timestep $t$. We have that the regret at timestep $t$ is
@@
3.5 Relationship with Littlestone dimension
We have so far shown that the Stackelberg-Littlestone dimension characterizes learnability in our setting. We next explore how that relates to multiclass learnability. While, as discussed, learning to predict the follower type from the context is sufficient to learn the leader’s optimal policy, in this section we prove that this is not a necessary condition. We start by showing that the Littlestone dimension upper bounds the Stackelberg Littlestone dimension. The proof of Lemma˜3.10 can be found in Section˜A.4.
Lemma 3.10.
For any hypothesis class $\pazocal{H}$ and Stackelberg game, $\pazocal{G}$, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})\leq\textnormal{Ldim}(\pazocal{H})$.
While it is known that any learner makes at least $\textnormal{Ldim}(\pazocal{H})$ mispredictions when learning a hypothesis class, $\pazocal{H}$, a careful reader might notice that this does not immediately show that the optimal multiclass learning algorithm (SOA) [22] achieves suboptimal regret in our setting. This is because mispredicting the follower type does not necessarily imply that the optimal strategy against the mispredicted and the true follower type do not yield the same utility for the leader. An illustrative example of this is our game construction in Theorem˜3.5, which shows that any learner can mispredict the follower type ad infinitum, yet the leader will achieve optimal cumulative utility regardless. In Theorem˜3.11, we prove that, indeed, there are instances where SOA will achieve smaller cumulative utility for the leader compared to our proposed algorithm. [^5] In fact, there are cases in which SOA yields strictly lower utility even against the algorithm that minimizes policy mispredictions. We prove this result in Lemma B.6.
Theorem 3.11.
There exists some hypothesis class $\pazocal{H}$ and a Stackelberg game, $\pazocal{G}$, such that running multiclass SOA on $\pazocal{H}$ yields suboptimal utility for the leader.
@@ 3.5 Relationship with Littlestone dimension
$h(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(2)}\;\;&\text{ if }z=2\\ f^{(1)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}\hskip 56.9055pth^{\prime}(z)=\begin{cases}f^{(3)}\;\;&\text{ if }z=1\\ f^{(1)}\;\;&\text{ if }z=2\\ f^{(2)}\;\;&\text{ if }z=3\\ f^{(2)}\;\;&\text{ if }z=4\\ \end{cases}$
Notice that the Littlestone dimension of this hypothesis class is at least 222. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜1.
Figure 1: Construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}$ and each edge corresponds to a follower type.
@@
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$. We will consider any possible Stackelberg Littlestone tree. First, note that without loss of generality we can consider trees that do not include the context $z=4$ since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, for $z=3$, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
So it suffices to consider contexts 111 and 222. In Figure˜2 we show the two shattered Stackelberg-Littlestone tree constructions with $z=1$ and $z=2$ as the root nodes respectively. We start with the tree in Figure˜2(a). For $z=2$, its not hard to show that $\inf_{\mathbf{x}}\max_{j\in\{1,3\}}r(\mathbf{x},2,f^{(j)})=\inf_{\mathbf{x}}\max_{j\in\{2,3\}}r(\mathbf{x},2,f^{(j)})=1/2$, achieved at $\mathbf{x}=(1/2,1/2)$. For the root node $z=1$, we have that $\inf_{\mathbf{x}}\max\{r(\mathbf{x},1,f^{(1)})+1/2,\;r(\mathbf{x},1,f^{(2)})+1/2,\;r(\mathbf{x},1,f^{(3)})\}=3/4$ achieved at $\mathbf{x}=(3/4,1/4)$. Therefore, the weight of the root node is 3/43/43 / 4. A similar analysis shows that the root node weight of the SL-tree in Figure˜2(b) is 3/43/43 / 4. Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals 3/43/43 / 4.
We consider the case, where the adversary present context 111 followed by context 222. We go through the SOA outputs per timestep below.
@@
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is 3/43/43 / 4 and we have shown that for the above sequence of contexts running SOA yields regret 222, this concludes our proof.
4 Distributional Learning in Structured Stackelberg Games
We now shift our focus from online learning to distributional learning. Consider an unknown, joint distribution over contexts and follower types, from which the leader receives samples drawn i.i.d. The leader’s objective is to design her strategy in such a way to maximize the utility she will get on a new sample drawn from the same distribution. We first study the simpler setting, without contexts, in which we give strong generalization guarantees by exploiting the structure of the leader’s payoff function. When contexts come into play in Section˜4.2, we prove necessary and sufficient conditions for PAC learnability.
4.1 Learning without Contexts
Suppose there is a fixed, unknown distribution, $\mathscr{D}$, over the set of possible follower types. We denote with $u(\mathbf{x},b_{f^{(i)}}(\mathbf{x}))$ the leader’s payoff when committing to a (possibly mixed) strategy $\mathbf{x}\in\Delta(\pazocal{A})$ against a follower of type $f^{(i)}$. We use the notation $u_{\mathscr{D}}(\mathbf{x})=\mathbb{E}_{f\sim\mathscr{D}}[u(\mathbf{x},b_{f}(\mathbf{x}))]$ and given a sample $S=\{f_{1},\dots,f_{m}\}$ of follower types, we write $u_{S}(\mathbf{x})=\frac{1}{m}\sum_{f\in S}u(\mathbf{x},b_{f}(\mathbf{x}))$.
Given such a sample, the goal of the leader is to choose a mixed strategy that fares well in expectation against any new sample from $\mathscr{D}$. Concretely, we want to choose $\hat{\mathbf{x}}\in\Delta(\pazocal{A})$ such that
$\mathbb{P}_{S}\left[\sup_{\mathbf{x}}u_{\mathscr{D}}(\mathbf{x})-u_{\mathscr{D}}(\hat{\mathbf{x}})>\epsilon\right]\leq\delta.$
Definition 4.1 (Generalization guarantee).
A generalization guarantee for the leader in a Stackelberg game, $\pazocal{G}$, is a function $\epsilon_{\pazocal{G}}:\mathbb{Z}_{\geq 1}\times(0,1)\to\mathbb{R}_{\geq 0}$ such that for any $\delta\in(0,1)$, any $m\in\mathbb{Z}_{\geq 1}$, and any distribution $\mathscr{D}$ over the follower types, with probability at least $1-\delta$ over the draw of a set $S\sim\mathscr{D}^{m}$ for any $\mathbf{x}\in\Delta(\pazocal{A})$, the difference between the average payoff of $\mathbf{x}$ over $S$ and the expected payoff of $\mathbf{x}$ over $\mathscr{D}$ is at most $\epsilon_{\pazocal{G}}(m,\delta)$:
$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[\exists\mathbf{x}\in\Delta(\pazocal{A})\;\;\text{such that}\;\;|u_{\mathscr{D}}(\mathbf{x})-u_{S}(\mathbf{x})|>\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta$
This implies a relationship between the utility of the leader’s strategy maximizing the average payoff over the training data to the expected utility of the optimal strategy. Formally, for $\hat{\mathbf{x}}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{S}(\mathbf{x})$ and $\mathbf{x}^{*}\in\arg\max_{\mathbf{x}\in\Delta(\pazocal{A})}u_{\mathscr{D}}(\mathbf{x})$
$\mathbb{P}_{S\sim\mathscr{D}^{m}}\left[|u_{\mathscr{D}}(\mathbf{x}^{*})-u_{S}(\hat{\mathbf{x}})|>2\epsilon_{\pazocal{G}}(m,\delta)\right]\leq\delta.$
We let $\pazocal{P}=\{u_{\mathbf{x}}:[K]\to\mathbb{R}\;\;\text{such that}\;\;u_{\mathbf{x}}(f)=u(\mathbf{x},b_{f}(\mathbf{x}))\}$ be the set of payoff functions for the leader. Each payoff function is parameterized by the leader’s mixed strategy, takes as input the follower’s type, and outputs the leader’s utility. We also define the dual class $\tilde{\pazocal{P}}=\{\tilde{u}_{f}:\Delta(\pazocal{A})\to\mathbb{R}\;\;\text{such that}\;\;\tilde{u}_{f}(\mathbf{x})=u(\mathbf{x},b_{f}(\mathbf{x}))\}$. Consider the following definition, introduced by Balcan et al. [9] and adapted to our setting.
Definition 4.2.
[…]
  • For any $f\in[K]$, there is a set $\mathscr{J}$ of $t$ hyperplanes such that for any connected component $\pazocal{X}^{\prime}=\pazocal{X}\setminus\mathscr{J}$, the dual function $h_{f}(\mathbf{x})$ is linear over $\pazocal{X}^{\prime}$.
Lemma 4.3.

### v3

@@ Appendix B Details and proofs from Section˜3
Observe that by construction, for any true hypothesis $h_{*}\in\pazocal{H}$ we have that the optimal policy $\pi_{h^{*}}$, will satisfy $\pi_{h^{*}}(\mathbf{z})=\pi_{*}^{(1)}(\mathbf{z})$ if $\mathbf{z}[2]<\frac{1}{2}$ and $\pi_{h^{*}}(\mathbf{z})=\pi_{*}^{(2)}(\mathbf{z})$ otherwise. Therefore, no matter which hypothesis is used to predict the follower’s type, the utility of the leader will be optimal. On the other hand, determining $h^{*}$ requires learning the exact cutoff values, $\gamma_{1}$ and $\gamma_{2}$. This corresponds to learning the class of linear thresholds, which has infinite Littlestone dimension (cf. Example 21.4 in [uml]). Therefore, $\textnormal{Ldim}(\pazocal{H})=\infty$. ∎
See 3.10
@@
We prove by induction on the depth $d$ of shattered trees that across $d$ timesteps the learner can be forced to incur regret at least as much as the weight of the tree’s root node. Since the SL dimension is defined as the supremum root node weight over all finite-depth trees, proving this suffices to show the desired claim. In case the supremum in the definition of $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ is not achieved by any finite tree, there will always be an additive error of $\epsilon$ that will get arbitrarily small as $d\to\infty$.
(Base Case) We consider a shattered tree of depth $1$. Then
@@
(Inductive step) We assume that for any tree of depth $d$ that is shattered by $\pazocal{H}$ under $\pazocal{G}$, the weight of the root node lower bounds the learner’s cumulative regret across $d$ timesteps. We now consider a tree $\mathcal{T}_{d+1}$ of depth $d+1$ that is shattered by $\pazocal{H}$ under Stackelberg game $\pazocal{G}$. We will prove that for such a tree, the adversary can guarantee regret at least $\rho_{\varnothing}$, where $\mathbf{z}_{\varnothing}$ is the root node of this tree. By Definition˜3.7, we know that $\rho_{\varnothing}=\inf_{\mathbf{x}\in\Delta(\pazocal{A})}\max_{i\in\pazocal{H}(\mathbf{z}_{\varnothing})}\biggl[r(\mathbf{z}_{\varnothing},\mathbf{x},f^{(i)})+\rho_{\varnothing i}\biggr]$. Let $\mathbf{x}_{*}\in\Delta(\pazocal{A})$ and $i_{*}\in[K]$ be the respective arginf and argmax values in the above expression.
@@
[…]
@@
See 3.11
@@
We will show that at each timestep $t$, if the learner incurred instantaneous regret $r_{t}=r(\mathbf{z}_{t},\mathbf{x}_{t},f_{t})$, then the the SL dimension must have decreased by at least $r_{t}$. In other words, $\textnormal{SLdim}_{\pazocal{G}}(V_{t+1})\leq\textnormal{SLdim}_{\pazocal{G}}(V_{t})-r_{t}$, where $V_{t}$ is the set of hypotheses consistent with the history up to timestep $t$. We have that the regret at timestep $t$ is
@@
See 3.12
@@
We assume for the sake of contradiction that there exists a hypothesis class with $\textnormal{Ldim}(\pazocal{H})=d$ and a Stackelberg game $\pazocal{G}$, such that $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})>d$. According to Definition˜B.5, this implies that there exists a shattered tree, whose minimum-weighted root-to-leaf path has weight strictly greater than $d$. This implies that all paths in the tree have cumulative weight strictly greater than $d$. Given that the weight of each edge in the tree is at most $1$, the length of the minimum path (and therefore of all other paths in the tree) must be of length at least $d+1$. This implies that there exists a tree of depth $d+1$ that is shattered by $\pazocal{H}$, which leads to a contradiction. ∎
See 3.14
Proof.
Consider $\pazocal{Z}=\{1,2,3,4\}$ and followers $\{f^{(1)},f^{(2)},f^{(3)}\}$. We define $\pazocal{H}=\{h_{\sigma}(z):=\mathbbm{1}\{z\neq 4\}\cdot f^{(\sigma(z))}+\mathbbm{1}\{z=4\}\cdot f^{(3)}\}_{\sigma\in\Sigma}\cup\{h,h^{\prime}\}$ , where $\Sigma$ is the set of all permutations of the set $\{1,2,3\}$ and $h,h^{\prime}$ are defined as follows.
@@
[…]
Notice that the Littlestone dimension of this hypothesis class is at least $2$. This follows from the $\pazocal{H}$-shattered tree construction shown in Figure˜4.
@@
$F_{1}$ $F_{2}$
$L_{1}$ $1,\mathbbm{1}_{\{z=3\}}/2$ $0,0$
$L_{2}$ $0,0$ $1,\mathbbm{1}_{\{z\neq 3\}}/2$
We now show that the $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=3/4$. We will consider any possible SL tree. First, note that without loss of generality we can consider trees that do not include the context $z=4$ since for $z=4,\;\pi_{*}^{(2)}(4)=\pi_{*}^{(3)}(4)$, where $f^{(2)}$ and $f^{(3)}$ are the only possible follower types predicted by hypotheses in $\pazocal{H}$ for this context. Similarly, for $z=3$, since the leader’s optimal strategy against all follower types is identical ($\pi_{*}^{(1)}(3)=\pi_{*}^{(2)}(3)=\pi_{*}^{(3)}(3)$).
So it suffices to consider contexts $1$ and $2$. In Figure˜5 we show the two shattered SL tree constructions with $z=1$ and $z=2$ as the root nodes respectively. We start with the tree in Figure˜5(a). For $z=2$, its not hard to show that $\inf_{\mathbf{x}}\max_{j\in\{1,3\}}r(\mathbf{x},2,f^{(j)})=\inf_{\mathbf{x}}\max_{j\in\{2,3\}}r(\mathbf{x},2,f^{(j)})=1/2$, achieved at $\mathbf{x}=(1/2,1/2)$. For the root node $z=1$, we have that $\inf_{\mathbf{x}}\max\{r(\mathbf{x},1,f^{(1)})+1/2,\;r(\mathbf{x},1,f^{(2)})+1/2,\;r(\mathbf{x},1,f^{(3)})\}=3/4$ achieved at $\mathbf{x}=(3/4,1/4)$. Therefore, the weight of the root node is $3/4$. A similar analysis shows that the root node weight of the SL-tree in Figure˜5(b) is $3/4$. Therefore, $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})$ (and also the worst-case cumulative regret) equals $3/4$.
We consider the case, where the adversary present context $1$ followed by context $2$. We go through the SOA outputs per timestep below.
  • ($t=1$) This node has three children corresponding to follower types $f^{(1)},f^{(2)},f^{(3)}$, with $f^{(1)}$ and $f^{(2)}$ leading to a full binary subtree of maximal depth 1 each. To see this, notice that (1) neither $h$ nor $h^{\prime}$ are consistent with the mapping of context $z=1$ to follower type $f^{(1)}$ or $f^{(2)}$, and (2) $\pazocal{H}\setminus\{h,h^{\prime}\}$ consists of all permutations mapping contexts $\{1,2,3\}$ to follower types $\{f^{(1)},f^{(2)},f^{(3)}\}$. Therefore, at most two of these contexts can exist in any root-to-leaf path of any $\pazocal{H}$-shattered tree. The child corresponding to follower type $f^{(3)}$ leads to a full binary subtree of maximal depth at least $2$, as shown in Figure˜4. So, the predicted follower type is $f_{1}=f^{(3)}$ and the strategy played is $x_{1}=\pi_{*}^{(3)}(1)$. However, the true follower type is $f_{1}^{*}=f^{(1)}$ (instantaneous regret equals 1).
This leads to SOA mispredicting the follower type and getting suboptimal utility in both rounds. Since the Stackelberg Littlestone dimension is $3/4$ and we have shown that for the above sequence of contexts running SOA yields regret $2$, this concludes our proof. ∎
Figure 4: Example construction of an $\pazocal{H}$-shattered Littlestone tree. Each node represents a context in $\pazocal{Z}=\{1,2,3,4\}$ and each edge corresponds to a follower type.
B.1 A technical subtlety
An important component of our analysis is that the leader must be able to compute the optimal strategy for a given context and follower type. For this, we state that for any context $\mathbf{z}\in\pazocal{Z}$ the leader incurs negligible regret by restricting herself to a subset $\mathcal{E}_{\mathbf{z}}\subseteq\Delta(\pazocal{A})$ which is finite and computable. To formally state our result we present the notion of a contextual best-response region, which was first introduced in [harrisregret] and which we repeat here for completeness.
Definition B.1 (Contextual Follower Best-Response Region).
Definition B.2 ($\delta$-approximate extreme points).
@@ B.1 A technical subtlety
Notice that for sufficiently small $\delta$ i.e. $\delta=\mathcal{O}(\frac{1}{T})$, we get that the additional regret of the leader by choosing actions over $\mathcal{E}^{(i)}$ rather than $X(f^{(i)},a_{f})$ is negligible. Therefore, we use the shorthand $\mathcal{E}^{(i)}:=\mathcal{E}^{(i)}(\delta)$ for the rest of the section. The following lemma from [harrisregret] formalizes this notion.
Lemma B.3.
For any sequence of followers $f_{1},\ldots f_{T}$ and any leader policy $\pi$, there exists a policy $\pi^{(\mathcal{E})}:\pazocal{Z}\rightarrow\cup_{\mathbf{z}\in\pazocal{Z}}\mathcal{E}_{\mathbf{z}}$ that, when given context $\mathbf{z}$, plays a mixed strategy in $\mathcal{E}_{\mathbf{z}}$ and guarantees that $\sum_{t=1}^{T}u(\mathbf{z}_{t},\pi(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi(\mathbf{z}_{t})))-u(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t}),b_{f_{t}}(\mathbf{z}_{t},\pi^{(\mathcal{E})}(\mathbf{z}_{t})))\leq 1$. Moreover, the same result holds in expectation over any distribution over follower types.
Given Lemma˜B.3 and for ease of notation, in this paper we assume that the leader will always set the parameter $\delta$, such that it does not affect our regret bounds.
B.2 Alternate definition of the Stackelberg Littlestone dimension
In this section we provide an alternate definition of the Stackelberg Littlestone dimension that can be defined for infinite trees. In Theorem˜B.6 we prove that for finite trees, this is equivalent to Definition˜3.9.
Definition B.4 ((Edge-weighted) SL tree).
An edge-weighted SL tree, $\pazocal{T}$, is a rooted tree, in which each internal node is labeled with an instance in $\pazocal{Z}$ and each edge corresponds to a label in $[K]$. For each node, $\mathbf{z}_{s}\in\pazocal{T}$ with children $\{\mathbf{z}_{si}\}_{i\in M\subseteq[K]}$ the outgoing edges have weights, $\{\nu_{s}^{(i)}\}_{i\in M}$, such that for every $\mathbf{x}\in\Delta(\pazocal{A})$ there exists $i\in M$ such that $r(\mathbf{z}_{s},\mathbf{x},f^{(i)})\geq\nu_{s}^{(i)}$. We say that the tree is shattered by a hypothesis class $\pazocal{H}$ under Stackelberg game $\pazocal{G}$ if for every path that traverses nodes $\mathbf{z}_{\varnothing},\dots\mathbf{z}_{s_{\leq d}}$, there exists $h\in\pazocal{H}$ such that the label of the edge in the path leaving $\mathbf{z}_{s\leq i}$ is $h(\mathbf{z}_{s_{\leq i}})$ for every $i$.
Definition B.5 ((Edge-weighted) SL dimension).
The SL dimension of a hypothesis class $\mathcal{H}$ with respect to a Stackelberg game, $\pazocal{G}$, is defined as $\textnormal{SLdim}_{\pazocal{G}}(\pazocal{H})=\sup\{\kappa\in\mathbb{R}_{\geq 0}:$ exists finite edge-weighted SL tree shattered by $\pazocal{H}$ under $\pazocal{G}$ where all root-to-leaf paths have cumulative weight at least $\kappa$}.
Theorem B.6.
For finite trees, Definition˜3.9 and Definition˜B.5 coincide.